A sample of an organic compound is found to have the following masses. Determine its
chemical formula.
Carbon 36.50 g
Hydrogen 9.19 g
Oxygen 24.31 g
1
Expert's answer
2018-12-26T11:11:37-0500

n=m/M

n(C)=36.50 g/ 12.01g/mol = 3.04 mol

n(H)=9.19 g/ 1.01g/mol =9.1 mol

n(O)= 24.31g/ 16.00 g/mol = 1.52 mol

n(C):n(H):n(O)=3.04 : 9.1 : 1.52

Divide each elemental ratio by the lowest number ( divide by 1.52)

n(C):n(H):n(O)=3.04/1.52 : 9.1/1.52 : 1.52/1.52 = 2:6:1

Chemical formula is C2H6O

Answer: C2H6O